# Generative Active Adaptation for Drifting and Imbalanced Network Intrusion Detection

### v1

@@ Abstract.
Machine learning has shown promise in network intrusion detection systems, yet its performance often degrades due to concept drift and imbalanced data. These challenges are compounded by the labor-intensive process of labeling network traffic, especially when dealing with evolving and rare attack types, which makes selecting the right data for adaptation difficult. To address these issues, we propose a generative active adaptation framework that minimizes labeling effort while enhancing model robustness. Our approach employs density-aware active sampling to identify the most informative samples for annotation and leverages deep generative models to synthesize diverse samples, thereby augmenting the training set and mitigating the effects of concept drift. We evaluate our end-to-end framework on both simulated IDS data and a real-world ISP dataset, demonstrating significant improvements in intrusion detection performance. Our method boosts the overall F1-score from 0.60 (without adaptation) to 0.86. Rare attacks such as Infiltration, Web Attack, and FTP-BruteForce, which originally achieve F1 scores of 0.001, 0.04, and 0.00, improve to 0.30, 0.50, and 0.71, respectively, with generative active adaptation in the CIC-IDS 2018 dataset. Our framework effectively enhances rare attack detection while reducing labeling costs, making it a scalable and adaptive solution for real-world intrusion detection.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Web protocol securityccs: Computing methodologies Batch learning
@@ 1. Introduction
Machine learning (ML) has significantly improved network intrusion detection systems (NIDS), surpassing traditional signature-based methods. Unlike manually curated signatures, ML-based NIDS can detect unknown attacks, especially with unsupervised techniques (Chou and Jiang, 2021). Various approaches — supervised (Holland et al., 2021; Liu et al., 2021; Xiao et al., 2022; Zhong et al., 2020; Shettar et al., 2021; Attia et al., 2020; Liu et al., 2024; Wan et al., 2024), unsupervised (Du et al., 2019; Liao and Li, 2022; Mirsky et al., 2018; Fu et al., 2021), and semi-supervised (Abdel-Basset et al., 2021; Jeong et al., 2021) — have been explored. Supervised models (e.g., CNN, MLP, XGBoost) achieve high accuracy but require large labeled datasets. Unsupervised methods (e.g., K-means, Kitsune (Mirsky et al., 2018)) use unlabeled data but often with lower accuracy. Semi-supervised techniques reduce labeling needs but may perform inconsistently on smaller datasets. Despite these challenges, ML- and DL-based NIDS have demonstrated over 95% detection accuracy in controlled tests (Sowmya and Mary Anita, 2023; Chou and Jiang, 2021).
@@
Despite their promise, current ML-driven NIDS frequently assume that training and testing data follow a static distribution. In reality, the network environment is subject to concept drift(Yang et al., 2021), where the behaviors of both benign and malicious traffic evolve over time. These changes arise from zero-day attacks, shifting user patterns, and instant updates to network services. In our research, a model trained on a well-known intrusion dataset (Sharafaldin et al., 2018) for one year saw its F1 score fall from 0.99 to 0.61 when tested on data from the following year, even though the types of attacks remained largely the same. Similar observations were made using the real-world UGR’16 (Maciá-Fernández et al., 2018) dataset, which contains anonymized NetFlow traces collected over four months in a Tier-3 ISP. This dataset integrates realistic attack scenarios with labeled traffic. When a classifier trained on one month of data (July) was tested on another month’s data (August), the F1 score dropped from 0.96 to 0.687, emphasizing the impact of temporal shifts on NIDS model performance.
@@
State of the Art. Traditional approaches to handle concept drift include retraining the model by combining new data with old data or retraining from scratch on the new data only (Pendlebury et al., 2019; Jan et al., 2020). However, these methods can be computationally expensive and risk catastrophic forgetting, where the model’s performance on previously learned data degrades as new information is introduced (French, 1999). Other advanced techniques for addressing concept drift in network intrusion detection rely on active learning to adapt models to evolving distributions (Chen et al., 2023a; Görnitz et al., 2009; Amalapuram et al., 2022). While numerous tools exist for collecting and capturing network data, labeling remains a significant challenge due to the time, expertise, and resources required. Active learning addresses this bottleneck by prioritizing the labeling of the most informative samples, enabling models to update effectively and maintain performance in dynamic environments. The adaptation of active learning methods in network intrusion detection mainly stems from different sampling strategies: uncertainty, diversity, density, or a combination of these. Uncertainty-based methods prioritize samples with the greatest predictive uncertainty, as seen in OWAD (Han et al., 2023). Diversity-based approaches, often combined with uncertainty sampling (e.g., ENiDrift (Wang, 2022)), seek to maximize the variability of labeled instances to improve generalization. Density-based methods, like CADE (Yang et al., 2021) and LEAF (Chen et al., 2023b), leverage embedding distances to identify structurally representative samples in feature space. We found that uncertainty-based methods struggle to select informative attack samples when majority of the benign samples are drifting, leading to biased models that miss rare attacks. Additionally, they require a high labeling budget, as many approaches still demand substantial annotations to sustain performance.
@@
Challenges. Building robust, learning-enabled NIDS is impeded by two major challenges: concept drift and class imbalance. (1) Network traffic patterns are highly dynamic, leading to significant concept drift. First, the evolution of normal traffic—driven by the proliferation of IoT devices, 5G/6G networks, and shifts in user behavior—continuously redefines baseline activity (Chou and Jiang, 2021). Second, attackers adapt their existing strategies, refining techniques to bypass detection (mit, 2022). Third, entirely new attack types can emerge suddenly, rendering previous models less effective (Layeghy and Portmann, 2023; Layeghy et al., 2023) (we illustrated this in Fig.1 (b) with new variants of DDoS attack in CIC 2018). Additionally, overlapping embeddings between benign and malicious traffic further blur the distinction between classes, complicating the task of accurate classification (Mirsky et al., 2018). (2) Class imbalance poses another critical challenge. NIDS datasets often exhibit a skewed distribution where benign samples vastly outnumber malicious ones, particularly for rare or emerging attack types (Chen et al., 2023c). This scarcity of minority class examples degrades the performance of ML models, as they become biased toward the majority class. Moreover, the low presence of attack samples hampers active adaptation; uncertainty-based active learning methods struggle to select informative samples when critical minority examples are underrepresented. This limitation not only reduces overall detection accuracy but also slows the system’s ability to adapt quickly to evolving threats.
@@
Our Work. In this work, we propose an-end-to-end framework, NetGuard , that employs a density-aware active learning strategy along with deep generative data augmentation to address concept drift and class imbalance between training (or source) and testing (or target) network intrusion detection datasets. We combine our own active learning method with data augmentation because rare attacks are underrepresented in NIDS datasets. Therefore, applying actively learned samples for retraining further exacerbates the problem of data scarcity. This is particularly problematic for training DNN-based network intrusion detection models, where reduced training data from active learning can lead to overfitting and inflated performance metrics for detection. To address this, there are two potential approaches; first, acquiring more samples during active learning which essentially contradicts our primary goal for optimizing labeling cost and second, synthetically generating more data points. The second approach takes the assumption that the synthesized dataset is representative of the original distribution. Therefore, NetGuard active learning and data augmentation collectively enhance the robustness and generalization of network intrusion detection models when the training and testing datasets differ due to temporal evolution or their origination from different network architecture or domains. Specifically, our contributions are as follows:
  • Density and Informativeness: Our algorithm employs a practical scenario with a labelled training dataset ($X_{L}$) and an unlabeled testing dataset ($X_{UL}$). The algorithm uses distribution-aware GMM scheme to identify samples in $X_{UL}$ that exhibit locally sparse density and differ significantly from the distribution of the training dataset $X_{L}$. This ensures that the selected samples are highly informative, representing areas of the feature space where the data distribution shifts, improving the overall sample quality for labeling.
@@
  • Diversity via Data Augmentation. To counter the challenges of class imbalance, our framework leverages deep generative models for data augmentation. By employing few-shot generation techniques, we synthesize diverse samples in underrepresented areas of the feature space. This augmentation not only creates a more balanced dataset but also improves the model’s ability to generalize to new and evolving attack types.
  • Evaluation. We validate our unified approach on multiple datasets: the popular CIC 2017/2018 datasets (Sharafaldin et al., 2018) and a real-world ISP network dataset, UGR’16 (Maciá-Fernández et al., 2018). The CIC-IDS 2017/2018 datasets originate from different network domains. CIC-IDS 2017 was collected on a smaller network architecture with 4 attack PCs, 10 victim PCs, 1 firewall, and 2 switches. In contrast, CIC-IDS 2018 was collected on a larger network with 10 attack PCs, 420 victim PCs, and a switch, leading to significant differences in traffic patterns and attack behaviors. The UGR16 dataset was collected over several months, and we evaluate NetGuard by using July as training data and August as testing data. Both CIC 2017/2018 and UGR’16 datasets consist of several classes, including rare attack types. This setup captures concept drift arising from both domain differences and temporal evolution with class imbalance, making it a suitable testbed for evaluating generative active adaptation strategies in NIDS.
@@
2. Background and Related Works
2.1. Active Learning
@@
Active Learning (AL) (Settles, 2009) reduces the labeling burden by selectively annotating the most informative samples, assuming an accurate labeling agent. Traditional uncertainty-based acquisition functions (Liu et al., 2023), such as entropy, margin sampling, and least confidence, struggle with imbalanced datasets, as minority class samples are often misclassified with high confidence and thus not selected for labeling (Aggarwal et al., 2020). Existing AL approaches for network intrusion detection (Stokes et al., 2008; Dang, 2020) focus on querying uncertain or rare/new attacks but fail to address concept drift, where prior class distributions may change. Other AL approaches, such as Coreset (Sener and Savarese, 2018), focus on data diversity to improve coverage across different class distributions. However, most of the AL methods either rely on model predictions or select samples without considering their relationship to previously seen data. This lack of comparison between new unlabeled data and previously trained data can lead to suboptimal sample selection.
2.2. Concept Drift Adaptations
In practice, most network intrusion detection systems suffer from narrow design scopes, focusing on limited aspects like concept drift or imbalanced data without addressing all key challenges simultaneously. Table 1 compares representative works against four features that are critical for real-world deployment: concept drift handling, imbalanced data mitigation, multi-class support, and labeling cost. For instance, CADE (Yang et al., 2021) and ENiDrift (Wang, 2022) both address concept drift, but only ENiDrift manages imbalanced data effectively; neither is designed to handle diverse multi-class scenarios or minimize labeling requirements. Similarly, (Yang et al., 2024) relies on self-supervised and weakly-supervised learning for concept drift adaptation in NIDS but continues to incur a large labeling cost. Additionally, it cannot improve the intrusion detection performance for underrepresented or rare attack types. LEAF (Chen et al., 2023b) focuses on explaining and mitigating concept drift in regression-based models for cellular networks but does not address supervised learning challenges such as labeling cost and class imbalance, which are critical in intrusion detection. In contrast, our proposed NetGuard framework aims to encompass all these dimensions—continuously adapting to drift, managing imbalance, supporting multi-class attacks, and reducing labeling overhead—thus offering a more holistic solution that aligns with real-world needs. By bridging these gaps, we seek to prevent attackers from exploiting the inherent weaknesses of specialized or narrowly focused NIDSs, ultimately providing stronger defense across dynamic and diverse network environments.
@@ 2.2. Concept Drift Adaptations
Features CADE (Yang et al., 2021) ENiDrift (Wang, 2022) ReCDA (Yang et al., 2024) LEAF (Chen et al., 2023b) NetGuard
@@
Sampling Density Uncert.+Divers. Density Density Density+Divers.
@@
2.3. Generative AI for Network Data Synthesis
Rare but critical attacks are underrepresented in network traffic, leading to high false negatives in learning-enabled NIDS. Additionally, publicly available network datasets are often obsolete, failing to capture evolving traffic patterns, while privacy concerns prevent ISPs and businesses from sharing real-world data. This scarcity limits the training of robust ML-based models for intrusion detection and traffic analysis. Recent advances in generative AI, including GANs (Lin et al., 2020; Yin et al., 2022), diffusion models (Jiang et al., 2024, 2023), GPT-based models (Zhang et al., 2024), state-space models (Chu et al., 2024), hybrid Transformer-GPT architectures (Solatorio and Dupriez, 2023), and VAEs (Xu et al., 2019), have shown promise in synthesizing high-fidelity network traces, improving model generalization and network analysis tasks. By learning from the underlying distribution of the original datasets, these generative models sample points from the learned distribution to expand the feature space and add diversity to the original training dataset.

### v2

@@ Abstract.
Machine learning has shown promise in network intrusion detection systems, yet its performance often degrades due to concept drift and imbalanced data. These challenges are compounded by the labor-intensive process of labeling network traffic, especially when dealing with evolving and rare attack types, which makes preparing the right data for adaptation difficult. To address these issues, we propose a generative active adaptation framework that minimizes labeling effort while enhancing model robustness. Our approach employs density-aware dataset prior selection to identify the most informative samples for annotation, and leverages deep generative models to conditionally synthesize diverse samples, thereby augmenting the training set and mitigating the effects of concept drift. We evaluate our end-to-end framework NetGuard on both simulated IDS data and a real-world ISP dataset, demonstrating significant improvements in intrusion detection performance. Our method boosts the overall F1-score from 0.60 (without adaptation) to 0.86. Rare attacks such as Infiltration, Web Attack, and FTP-BruteForce, which originally achieved F1 scores of 0.001, 0.04, and 0.00, improve to 0.30, 0.50, and 0.71, respectively, with generative active adaptation in the CIC-IDS 2018 dataset. Our framework effectively enhances rare attack detection while reducing labeling costs, making it a scalable and practical solution for intrusion detection.
@@
copyright: acmlicensedconference: ; ; isbn: XXXccs: Security and privacy Web protocol securityccs: Computing methodologies Batch learning
@@ 1. Introduction
Machine learning (ML) has significantly improved network intrusion detection systems (NIDS), surpassing traditional signature-based methods. Unlike manually curated signatures, ML-based NIDS can detect unknown attacks, especially with unsupervised techniques (Chou and Jiang, 2021). Various approaches — supervised (Holland et al., 2021; Liu et al., 2021; Xiao et al., 2022; Zhong et al., 2020; Shettar et al., 2021; Attia et al., 2020a; Liu et al., 2024; Wan et al., 2024), unsupervised (Du et al., 2019; Liao and Li, 2022; Mirsky et al., 2018; Fu et al., 2021), and semi-supervised (Abdel-Basset et al., 2021; Jeong et al., 2021) — have been explored. Supervised models (e.g., CNN, MLP, XGBoost) achieve high accuracy but require large labeled datasets. Unsupervised methods (e.g., K-means, Kitsune (Mirsky et al., 2018)) use unlabeled data but often with lower accuracy. Semi-supervised techniques reduce labeling needs but may perform inconsistently on smaller datasets. Despite these challenges, ML- and DL-based NIDS have demonstrated over 95% detection accuracy in controlled tests (Sowmya and Mary Anita, 2023; Chou and Jiang, 2021).
@@
Despite their promise, current ML-driven NIDS frequently assume that training and testing data follow a static distribution. In reality, the network environment is subject to concept drift(Yang et al., 2021), where the statistical behaviors of both benign and malicious traffic evolve, thus the models trained on the original data fail to capture the changes. These changes arise from zero-day attacks, shifting user patterns, and instant updates to network services. In our research, a model trained on a well-known intrusion dataset (Sharafaldin et al., 2018) for one year saw its F1 score fall from 0.99 to 0.61 when tested on data from the following year, even though the types of attacks remained largely the same. Similar observations were made using the real-world UGR’16 (Maciá-Fernández et al., 2018) dataset, which contains anonymized NetFlow traces collected over four months in a Tier-3 ISP. This dataset integrates realistic attack scenarios with labeled traffic. When a classifier trained on one month of data (July) was tested on another month’s data (August), the F1 score dropped from 0.96 to 0.687, emphasizing the impact of temporal shifts on the NIDS model.
@@
State of the Art. Traditional approaches to handle concept drift include retraining the model by combining new data with old data or retraining from scratch on the new data only (Pendlebury et al., 2019; Jan et al., 2020). However, these methods can be computationally expensive and risk catastrophic forgetting, where the model’s performance on previously learned data degrades as new information is introduced (French, 1999). Other advanced techniques for addressing concept drift in network intrusion detection rely on active learning to adapt models to evolving distributions (Chen et al., 2023a; Görnitz et al., 2009; Amalapuram et al., 2022a). While numerous tools exist for collecting and capturing network data, labeling remains a significant challenge due to the time, expertise, and resources required. Active learning addresses this bottleneck by prioritizing the labeling of the most informative samples, enabling models to update effectively and maintain performance in dynamic environments. The adaptation of active learning methods in network intrusion detection mainly stems from different sampling strategies: uncertainty, diversity, density, or a combination of these. Uncertainty-based methods prioritize samples with the greatest predictive uncertainty, as seen in OWAD (Han et al., 2023). Diversity-based approaches, often combined with uncertainty sampling (e.g., ENiDrift (Wang, 2022)), seek to maximize the variability of labeled instances to improve generalization. Density-based methods, like CADE (Yang et al., 2021) and LEAF (Chen et al., 2023b), leverage embedding distances to identify structurally representative samples in feature space. We found that uncertainty-based methods struggle to select informative attack samples when the majority of the benign samples are drifting, leading to biased models that miss rare attacks. Additionally, they require a high labeling budget, as many approaches still demand substantial annotations to sustain performance.
@@
Challenges. Building robust, learning-enabled NIDS is impeded by two major challenges: Concept Drift and Class Imbalance. (1) Network traffic patterns are highly dynamic, leading to significant concept drift. First, the evolution of normal traffic—driven by the proliferation of IoT devices, 5G/6G networks, and shifts in user behavior—continuously redefines benign activities (Chou and Jiang, 2021). Second, attackers adapt their existing strategies, refining techniques to bypass detection (mit, 2022). Third, entirely new attack types can emerge suddenly, rendering previous models less effective (Layeghy and Portmann, 2023; Layeghy et al., 2023) (we illustrated this in Fig.1 (b) with new variants of DDoS attack in CIC 2018). Additionally, overlapping embeddings between benign and malicious traffic further blur the distinction between classes, complicating the task of accurate classification (Mirsky et al., 2018). (2) Class imbalance poses another critical challenge. NIDS datasets often exhibit a skewed distribution where benign samples vastly outnumber malicious ones, particularly for rare or emerging attack types (Chen et al., 2023c). This scarcity of minority class examples degrades the performance of ML models, as they become biased toward the majority class. Moreover, the low presence of attack samples hampers active adaptation; uncertainty-based active learning methods struggle to select informative samples when critical minority examples are underrepresented. This limitation not only reduces overall detection accuracy but also slows the system’s ability to adapt quickly to evolving threats.
@@
Our Contributions. We propose an end-to-end framework, called NetGuard , that employs a density-aware dataset prior selection strategy along with conditional generative data augmentation to address concept drift and class imbalance between training (or source) and testing (or target) network intrusion detection datasets. The drifted and rare attacks are underrepresented in NIDS datasets. Naïvely applying actively learned samples for retraining further exacerbates the problem of data scarcity. This is particularly problematic for training NIDS models, where reduced training data can lead to overfitting and inflated performance metrics for detection. To address this, there are two potential approaches: first, acquiring more samples, which essentially contradicts our goal for optimizing labeling cost, and second, synthetically generating more data points. The second approach takes the assumption that the synthesized dataset is representative of the original distribution. Thus, NetGuard identifies a dataset prior and conditionally synthesizes more diverse data, thus enhancing the robustness and generalization of NIDS models when the training and testing datasets differ due to temporal evolution or their origination from different network architectures or domains. We make the following contributions. A. Density and Informativeness. Our algorithm employs a practical scenario with a labeled training dataset ($X_{L}$) and an unlabeled testing dataset ($X_{UL}$). The algorithm uses a distribution-aware GMM scheme to identify samples in $X_{UL}$ that exhibit locally sparse density and differ significantly from the distribution of the training dataset $X_{L}$. This ensures that the selected dataset priors are highly informative, representing areas of the feature space where the data distribution shifts, improving the sample quality for labeling.
@@
B. Diversity via Data Augmentation. To counter the challenges of class imbalance, our framework leverages deep generative models for conditional data augmentation on the dataset priors. By employing few-shot generation techniques, we synthesize diverse samples in underrepresented areas of the feature space. This augmentation not only creates a more balanced dataset but also improves the model’s ability to generalize to new and evolving attack types.
2. Background
@@
2.1. Related Works
Concept Drift Adaptations. Concept drift is a key challenge in real-world NIDS due to evolving protocols, devices, user behavior, and new attack vectors. Foundational Surveys (Gama et al., 2014; Lu et al., 2019; Webb et al., 2018, 2016) emphasize its prevalence and negative impact on detection accuracy. However, most existing NIDS solutions handle either drift or class imbalance, not both. Table 1 compares representative works against four features that are critical for real-world deployment: concept drift handling, imbalanced data mitigation, multi-class support, and labeling cost. For instance, CADE (Yang et al., 2021) and ENiDrift (Wang, 2022) both address concept drift, but only ENiDrift manages imbalanced data effectively; neither is designed to handle multi-class scenarios or minimize labeling requirements. ReCDA (Yang et al., 2024) adapts to drift using self-supervision but still incurs high labeling cost and struggles with rare attacks. LEAF (Chen et al., 2023b) focuses on drift in regression tasks for cellular networks, but lacks support for supervised, imbalanced, multi-class intrusion detection. In contrast, NetGuard addresses all four dimensions—adapting to drift, managing imbalance, supporting multi-class attacks, and reducing labeling overhead—making it more practical for real-world NIDS. In contrast, NetGuard framework aims to encompass all these dimensions—continuously adapting to drift, managing imbalance, supporting multi-class attacks, and reducing labeling overhead—thus offering a solution that aligns with real-world needs. Recent zero-shot and meta-learning approaches (Wang et al., 2022; Li et al., 2023a, b; Zhang et al., 2024b) adapt quickly in unsupervised domains like time series or images, but they are not suited for supervised NIDS. These methods detect deviations from “normal” without labels, but fail when class labels persist under shift. Their reliance on meta-distributions also limits their use in open-world scenarios where novel attacks lack prior structure.
@@ 2.1. Related Works
Active Learning. Active Learning (AL) (Settles, 2009) reduces the labeling burden by selectively annotating the most informative samples, assuming an accurate labeling agent. Traditional uncertainty-based acquisition functions (Liu et al., 2023), such as entropy, margin sampling, and least confidence, struggle with imbalanced datasets, as minority class samples are often misclassified with high confidence and thus not selected for labeling (Aggarwal et al., 2020). Existing AL approaches for network intrusion detection (Stokes et al., 2008; Dang, 2020) focus on querying uncertain or rare/new attacks but fail to address concept drift, where prior class distributions may change. Other AL approaches, such as Coreset (Sener and Savarese, 2018), focus on data diversity to improve coverage across different class distributions. However, most of the AL methods either rely on model predictions or select samples without considering their relationship to previously seen data. This lack of comparison between new unlabeled data and previously trained data can lead to suboptimal sample selection. Recent methods address fast or zero-shot anomaly detection under distributional shift, primarily in unsupervised domains such as multivariate time series or image data  (Wang et al., 2022; Li et al., 2023a, b; Zhang et al., 2024b). These approaches leverage meta-learning or prototype transfer for hypernetwork tuning to enable rapid adaptation, but they are not tailored for supervised NIDS with tasks with class-labeled and imbalanced attack classes. While effective at detecting deviations from the ”normal” class without labels, types. These methods can identify samples that deviate from ”normal” without requiring class labels, but they cannot prioritize high-impact samples when class labels remain fixed under distributional shift. They also rely on access to meta-distributions or related tasks during training (Wang et al., 2022; Li et al., 2023a, b), which enables few-shot generalization but limits applicability in open-world settings, real-time network environments where novel unseen attacks emerge without prior task structure.
Generative AI for Network Data Synthesis Rare but critical attacks are underrepresented in network traffic, leading to high false negatives in learning-enabled NIDS. Additionally, publicly available network datasets are often obsolete, failing to capture evolving traffic patterns, while privacy concerns prevent ISPs and businesses from sharing real-world data. This scarcity limits the training of robust ML-based models for intrusion detection and traffic analysis. Recent advances in generative AI, including GANs (Lin et al., 2020; Yin et al., 2022), diffusion models (Jiang et al., 2024, 2023), GPT-based models (Zhang et al., 2024a), state-space models (Chu et al., 2024), hybrid Transformer-GPT architectures (Solatorio and Dupriez, 2023), and VAEs (Apellaniz et al., 2024; Xu et al., 2019), have shown promise in synthesizing high-fidelity network traces to expand the feature space and improve model generalization for network analysis. However, most of these methods focus on zero-shot augmentation, which requires domain expertise or class-conditional generation using all available samples, without accounting for evolving distributions during test time. Other augmentation works, such as RAPIER (Qing et al., 2024) addresses label noise via GAN-based augmentation under static distributions, a problem orthogonal to our framework where we assume labels are correct under dynamic data distributions. To the best of our knowledge, prior works in building robust NIDS address either concept drift within a single dataset (Yang et al., 2021, 2024) or apply generic data augmentation for class imbalance (Yin et al., 2022). In contrast, we propose a novel approach that integrates state-of-the-art generative models with conditional generation, guided by drift-aware prior selection. This enables targeted synthesis of minority-class traffic under evolving distributions, offering a practical solution for adaptive NIDS in real-world settings. Our work is among the first to examine cross-domain drift using feature-aligned datasets (CIC-IDS 2017 and 2018), offering a realistic test of model robustness across deployments. For datasets like UGR’16 with different features, we instead evaluate temporal drift within a single domain.
@@
Features CADE (Yang et al., 2021) ENiDrift (Wang, 2022) ReCDA (Yang et al., 2024) LEAF (Chen et al., 2023b) CLUE (Prabhu et al., 2021) NetGuard
@@
Sampling Density Uncert.+Div. Density Density Uncert.+Div. Dens.+Div.